# Quantum-Enhanced Power Flow and Optimal Power Flow based on Combinatorial Reformulation

### v2

@@ III-B Combinatorial Optimal Power Flow Formulation
where $a_{i,j}$ are binary variables representing the discretized values of the slack variable $s_{i}$. By substituting expression (24) into 23, the original inequality constraints can be transformed into an equivalent QUBO formulation:
@@
IV Quantum and Digital Annealing
Quantum annealing is an analog computing paradigm designed to solve unconstrained optimization problems formulated as an Ising model [17]. The Ising Hamiltonian reads:
@@ IV Quantum and Digital Annealing
where $J_{ij}$ represents interaction coefficients between spins, and $h_{i}$ represents an external field, which acts as a linear bias that influences the tendency of $s_{i}$ toward +1+1 and 1-1; $\langle i,j\rangle$ represents all connections between $i,j\in\{1,\dots,n\}$ such that $i<j$. The system follows the dynamics of the time-dependent Schrödinger equation, starting from a highly delocalized quantum state and evolves according to a time-varying Hamiltonian. This evolution is designed such that, under ideal conditions, the system settles into the ground state of the final Hamiltonian, which corresponds to the optimal solution [18]. The QUBO formulations derived in Section III can be converted into Ising models by setting $x_{i}=(s_{i}-1)/2$ and regrouping the coefficients into $J_{ij}$ and $h_{i}$.
@@
Quantum and Digital Annealer (DA) are specialized hardware architectures designed to minimize the energy of quadratic functions over binary variables using this optimization approach. These architectures are referred to as Ising machines [19, 20]. D-Wave’s Advantage™ system [^1] https://www.dwavesys.com is one of the most well-known implementations of quantum annealing, featuring over 5,000 qubits and 35,000 couplers. In this system, a positive weight $J_{ij}>0$ encourages opposite spin orientations (ferromagnetic interaction), while a negative weight $J_{ij}<0$ favors aligned spins (antiferromagnetic interaction). Similarly, $h_{i}>0$ implies that the spin at site $i$ favors line up in the positive direction, whereas $h_{i}<0$ implies the opposite. Practical challenges arise due to the limited qubit connectivity, necessitating minor embedding to map the problem graph onto the hardware topology, an NP-hard problem. In addition, the range of $J$ and $h$ values is restricted, thus leading to a potential loss of precision in problem encoding.
@@
To overcome these challenges, Fujitsu introduced the DA, initially designed as an application-specific complementary metal-oxide semiconductor (CMOS) hardware [^2] https://www.fujitsu.com/global/services/business-services/digital-annealer, which emulates simulated annealing. It supports massively parallel execution of Markov Chain Monte Carlo (MCMC) processes and offers significantly higher qubit connectivity than quantum annealing. Fujitsu’s Digital Annealer Unit (DA) can handle up to 100,000 fully connected binary variables with 64-bit precision. Unlike quantum annealing, which requires, depending on the utilized qubit technology, cryogenic temperatures, digital annealing operates at room temperature, making it suitable for broader deployment, including edge computing applications. Each run of DA begins with a randomly generated binary configuration and iteratively explores the solution space to identify a state that yields minimal energy. Another key advantage of digital annealing over quantum annealing is its ability to natively accept problems in QUBO form. Building on simulated annealing, the second generation DA uses parallel tempering, a method that simulates multiple solution paths at different temperature levels and enables configuration swaps between them to improve exploration and prevent stagnation in local minima.
@@
A significant limitation of both quantum and digital annealing, however, is their restriction to binary quadratic cost functions. Many real-world problems, such as OPF, require handling higher-order polynomial cost functions and inequality constraints. Recent advancements in DA introduce support for higher-order terms and constrained optimization. DA achieves this by employing different operations to reduce the computational cost of multiplications and applying replica exchange Monte Carlo methods to enhance search efficiency. These improvements significantly expand the range of applications beyond traditional QUBO problems, including applications in large-scale power system [15]. The software implementation of [15] is available as ”Quantum-Inspired Integrated Optimization” (QIIO) [^3] https://en-portal.research.global.fujitsu.com/kozuchi, name for the latest DA prototype.
IV-A Higher-Order Terms Handling
@@
Expanding the full expressions of 15 for power flow (PF) and 29 for optimal power flow (OPF) results in a fourth-degree polynomial involving $\mathbf{x}$. QIIO supports native optimization over higher-order polynomials, but DAv3, QA, and HA are limited to solving problems with at most quadratic terms. Consequently, higher-order interactions must be reformulated into equivalent quadratic representations to be compatible with these solvers. One standard method for reducing cubic terms is to introduce auxiliary variables, such as $z_{ij}=x_{i}x_{j}$ allowing triplet products to be reformulated as:
@@
$x_{i}x_{j}x_{k}x_{l}=\min_{z_{ij},z_{kl}}\left\{\lambda P(x_{i},x_{j};z_{ij})+\lambda P(x_{k},x_{l};z_{kl})\right\}$ (33)
@@
For an in-depth review of quadratization approaches, see [21]. To construct the QUBO formulations corresponding to 15 for PF and 29 for OPF but also reducing higher-order terms, we employ the PyQUBO package when targeting QA and HA solvers, and the Python package DADK when targeting DAv3. After solving the minimization problem, the resulting bitstring is used to update $\mu_{i}$ and $\omega_{i}$ according to 12. Note that we neglect the auxiliary variables $z_{ij}$ for the bus complex voltage updates.
@@
Our experiments reveal discrepancies between the PyQUBO and DADK implementations, particularly in the number of auxiliary variables generated when encoding identical problem sizes. In general, DADK tends to produce more compact formulations with fewer auxiliary variables. We rigorously verify that the order reduction preserves the optimality of solutions and maintains consistency in variable assignments corresponding to the minimum-energy configurations.
V Adiabatic Quantum Algorithm
@@
Iterative schemes are introduced to ensure that AQPF and AQOPF use the unique computational properties of quantum/digital annealers to perform combinatorial PF and OPF, respectively, as presented in Algorithm 1. Initially, the active power generation vector $\mathbf{P}^{G}$, the demand vectors $\mathbf{P}^{D}$ and $\mathbf{Q}^{D}$, along with the power system admittance matrix $\mathbf{Y}$, are defined (lines 1-4). Next, the voltage real and imaginary parts, $\mathbf{\mu}$ and $\mathbf{\omega}$, as well as their corresponding adjustments, $\Delta{\mu}$ and $\Delta{\omega}$, are initialized by values provided by user (refer to lines 5–8). These parameters are then used to compute the starting values of the active and reactive power vectors, $\mathbf{P}$ and $\mathbf{Q}$ (see lines 9–10). Note that $P_{1}$ and $Q_{1}$, associated with the slack bus, are excluded to align with PF analysis.
@@
On line 11, the Hamiltonian $H_{\text{obj}}(\cdot)$ is evaluated at the binary vector $\mathbf{x}=[x_{0,0}^{\mu},\dots,x_{n,1}^{\omega}]\in\{0\}^{4n}$ (line 11) using the initial voltage estimations to propose a candidate solution. In line 12, the residual limit, defined as $\epsilon$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, is initialized, followed by resetting the iteration counter ‘it’ to zero prior to entering the iterative loop at line 13. After solving the minimization problem 15 (line 15), the binary solution vector $\mathbf{x}=[x_{0,0}^{\mu},\dots,x_{n,1}^{\omega}]\in\{0,1\}^{4n}$ derived from the QUBO formulation is employed to revise $\mu_{i}$ and $\omega_{i}$ as specified in line 16 by 12. It is important to note that during the voltage update process, any auxiliary variables introduced through the quadratic reformulation of higher-order terms are intentionally omitted. When the computed $H_{\text{obj}}(\mathbf{x})$ falls below the user-defined threshold $\epsilon$, the resulting complex voltages $\mathbf{\mu}+j\mathbf{\omega}$ are considered as the final solution. If not, the base values, $\mu_{i}^{0}$ and $\omega_{i}^{0}$, are reassigned from the current $\mu_{i}$ and $\omega_{i}$ (line 20), and the optimization routine for 15 is repeated until convergence is achieved.
@@
For AQOPF, Algorithm 1 should be modified to align with the OPF problem principle. First of all, line 1 is not needed, as the generator set points are obtained by solving the OPF problem. Moreover, the active and reactive power vectors, $\mathbf{P}$ and $\mathbf{Q}$, are computed only for load buses in lines 9, 10, 17, and 18. Finally, the problem Hamiltonian in lines 11 and 18 should be replaced with 29.
@@
9:Calculate $\mathbf{P}=[P_{2},P_{3},\dots,P_{N}]$ using 13
10:Calculate $\mathbf{Q}=[Q_{2},Q_{3},\dots,Q_{N}]$ using 14
11:Calculate Hamiltonian $H_{\text{obj}}(\mathbf{x})$ using 15
@@
16:  Update $\mathbf{\mu}$ and $\mathbf{\omega}$ using 12
17:  Recalculate $\mathbf{P}$ using 13
18:  Recalculate $\mathbf{Q}$ using 14
19:  Recalculate $H_{\text{obj}}(\mathbf{x})$ using 15
@@
V-A Delta Update
The iterative method may exhibit slow convergence when $\Delta\mu_{i}$ and $\Delta\omega_{i}$ are set too low, or if $\mu_{i}^{0}$ and $\omega_{i}^{0}$ are significantly different from the values that fulfill 15 and 29. A possible remedy is to employ an adaptive strategy for $\Delta\mu_{i}$ and $\Delta\omega_{i}$ where their values are dynamically adjusted per iteration per bus $i$ to enhance convergence behavior. In this perspective, the values for $\Delta\mu_{i}$ and $\Delta\omega_{i}$ are a function of the iteration counter ‘it’, as given in:
@@ V-A Delta Update
where $\underline{\Delta\mu_{i}}=5\times 10^{-4}$ and $\overline{\Delta\mu_{i}}=4\times 10^{-2}$ are lower and upper bounds of $\Delta\mu_{i}$, respectively; Similarly, $\underline{\Delta\omega_{i}}=1\times 10^{-4}$ and $\overline{\Delta\omega_{i}}=2\times 10^{-2}$ bound $\Delta\omega_{i}$; $e$ is the exponential function; r=0.05r=-0.05 is the rate of decay. In each iteration, $\Delta\mathbf{\mu}$ and $\Delta\mathbf{\omega}$ are updated adaptively and individually; we save two sets of binary variables $\mathbf{x}^{(\text{it}-1)}$ and $\mathbf{x}^{(\text{it}-2)}\in\{0,1\}^{4n}$, which represent the bitstrings for the previous and second previous iterations, respectively. We then use these bitstrings to update $\Delta\mu_{i}$ and $\Delta\omega_{i}$ for each bus $i$. For example, $x_{i,0}^{\mu(\text{it}-2)}=1\land x_{i,1}^{\mu(\text{it}-2)}=0$, $x_{i,0}^{\mu(\text{it}-1)}=1\,\land\,x_{i,1}^{\mu(\text{it}-1)}=0$, and $x_{i,0}^{\mu}=0\land x_{i,1}^{\mu}=1$ indicates that at iteration $\text{it}-2$, the value of $\mu_{i}$ increased, followed by an increase at iteration $\text{it}-1$ and then a decrease at the current iteration ‘it’. This pattern suggests that to prevent $\mu_{i}$ from fluctuations and promote convergence, we should reduce $\Delta\mu_{i}$. The full details are given in Algorithm 2. We also employ a similar approach to update $\Delta\mathbf{\omega}$.
@@
VI Results
Four annealers are employed: D-Wave’s Advantage™  system (QA), its hybrid quantum-classical solver (HA), as well as the third-generation Digital Annealer (DAv3) and Quantum-Inspired Integrated Optimization software (QIIO) developed by Fujitsu. The number of readout samples per execution is fixed for all solvers, with values between 2,000 and 100,000, determined by the test case size. A time limit of 10 seconds is assigned to HA. Minor embedding is employed for QA, with the chain strength selected to minimize chain breaks while preserving the relative importance of the original QUBO coefficients. The embedding target graph exhibits sparsity, mirroring the inherently low connectivity structure typical of power systems. For DAv3 and QIIO, the runtime configuration includes a 10-second time limit, 64-bit scaling precision, and a total optimization timeout of 3,600 seconds. Table I summarizes the capability of the solvers to converge to a solution within the residual tolerance of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$. For QA and HA, convergence indicates that an embedding is successfully found, and the solution is obtained within the accepted residual tolerance. For DAv3 and QIIO, convergence implies that the solution is obtained within the same tolerance.
According to Table I, QIIO is capable of handling all test case sizes, including the 1354-bus test case. DAv3, on the other hand, is limited to test cases up to the 89-bus test case due to the constraints on the number of available binary variables. QA and HA are even more restricted and are applicable only for small test cases. The difference between DAv3 and QIIO lies in their hardware configurations: DAv3 provides a fully connected block of approximately 8,192 binary variables, while QIIO offers multiple fully connected blocks capable of managing up to 100,000 binary variables. This expanded capacity makes QIIO suitable for large-scale problems. Therefore, all further experiments in this paper are conducted using QIIO. Note that the experiments include test cases ranging from 4- to 1354-bus test cases from the Power System Test Cases library [^4] https://pandapower.readthedocs.io/en/v2.2.2. The results for the 4-, 9-, and 14-bus test cases are presented in [16] and the results for 5-, 6-, and 15-bus test cases are presented in [22].
TABLE I: Performance comparison of different solvers under a user-defined residual of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$.

### v3

@@ 3.2 Combinatorial Optimal Power Flow Formulation
where $a_{i,j}$ are binary variables representing the discretized values of the slack variable $s_{i}$. By substituting expression (24) into ˜23, the original inequality constraints can be transformed into an equivalent QUBO formulation:
@@
4 Quantum and Digital Annealing
Quantum annealing is an analog computing paradigm designed to solve unconstrained optimization problems formulated as an Ising models Goto2019 . The Ising Hamiltonian reads:
@@ 4 Quantum and Digital Annealing
where $J_{ij}$ represents interaction coefficients between spins, and $h_{i}$ represents an external field, which acts as a linear bias that influences the tendency of $s_{i}$ toward +1+1 and 1-1; $\langle i,j\rangle$ represents all connections between $i,j\in\{1,\dots,n\}$ such that $i<j$. The system follows the dynamics of the time-dependent Schrödinger equation, starting from a highly delocalized quantum state and evolves according to a time-varying Hamiltonian. This evolution is designed such that, under ideal conditions, the system settles into the ground state of the final Hamiltonian, which corresponds to the optimal solution Lucas2014 . The QUBO formulations derived in Section˜3 can be converted into Ising models by setting $x_{i}=(s_{i}-1)/2$ and regrouping the coefficients into $J_{ij}$ and $h_{i}$.
@@
Quantum and Digital Annealer (DA) are specialized hardware architectures designed to minimize the energy of quadratic functions over binary variables using this optimization approach. These architectures are referred to as Ising machines Bunyk2014 ; Goto2021 . D-Wave’s Advantage™ system [^1] https://www.dwavesys.com is one of the most well-known implementations of quantum annealing, featuring over 5,000 qubits and 35,000 couplers. In this system, a positive weight $J_{ij}>0$ encourages opposite spin orientations (ferromagnetic interaction), while a negative weight $J_{ij}<0$ favors aligned spins (antiferromagnetic interaction). Similarly, $h_{i}>0$ implies that the spin at site $i$ favors lining up in the positive direction, whereas $h_{i}<0$ implies the opposite. Practical challenges arise due to the limited qubit connectivity, necessitating minor embedding to map the problem graph onto the hardware topology, an NP-hard problem. In addition, the range of $J$ and $h$ values is restricted, thus leading to a potential loss of precision in problem encoding.
@@
To overcome these challenges, Fujitsu introduced the DA, an application-specific complementary metal-oxide semiconductor (CMOS) hardware [^2] https://www.fujitsu.com/global/services/business-services/digital-annealer that emulates simulated annealing. It supports massively parallel execution of Markov Chain Monte Carlo (MCMC) processes and offers significantly higher qubit connectivity than quantum annealing. Fujitsu’s DA Unit can handle up to 100,000 fully connected binary variables with 64-bit precision. Unlike quantum annealing, which, depending on the qubit technology used, requires cryogenic temperatures, digital annealing operates at room temperature, making it suitable for broader deployment, including edge computing applications. Each run of DA begins with a randomly generated binary configuration and iteratively explores the solution space to identify a state that yields minimal energy. Another key advantage of digital annealing over quantum annealing is its native support for QUBO problems Inoue2024 . Building on simulated annealing, the second-generation DA uses parallel tempering. This method simulates multiple solution paths at different temperature levels and enables configuration swaps between them to improve exploration and prevent stagnation in local minima.
@@
A significant limitation of both quantum and digital annealing, however, is their restriction to binary quadratic cost functions. Many real-world problems, such as OPF, require handling higher-order polynomial cost functions and inequality constraints. Recent advancements in DA introduce support for higher-order terms and constrained optimization. DA achieves this by employing various operations to reduce the computational cost of multiplications and by applying replica-exchange Monte Carlo methods to enhance search efficiency. These improvements significantly expand the range of applications beyond traditional QUBO problems, including large-scale power systems. The software implementation of Yin2023 is available as QIIO [^3] https://en-portal.research.global.fujitsu.com/kozuchi, the name for the latest DA prototype.
4.1 Higher-Order Terms Handling
@@
Expanding the full expressions of ˜15 for PF and ˜29 for OPF results in a fourth-degree polynomial involving $\mathbf{x}$. QIIO supports native optimization over higher-order polynomials, but DAv3, QA, and HA are limited to quadratic terms. Consequently, higher-order interactions must be reformulated into equivalent quadratic representations to be compatible with these solvers. One standard method for reducing cubic terms is to introduce auxiliary variables, such as $z_{ij}=x_{i}x_{j}$, allowing triplet products to be reformulated as:
@@
$x_{i}x_{j}x_{k}x_{l}=\min_{z_{ij},z_{kl}}\left\{\lambda P(x_{i},x_{j};z_{ij})+\lambda P(x_{k},x_{l};z_{kl})\right\}.$ (33)
@@
For an in-depth review of quadratization approaches, see Dattani2019 . To construct the QUBO formulations corresponding to ˜15 for PF and ˜29 for OPF, while also reducing higher-order terms, we employ the PyQUBO package for QA and HA solvers, and the Python package DADK for DAv3. After solving the minimization problem, the resulting bitstring is used to update $\mu_{i}$ and $\omega_{i}$ according to ˜12. Note that we neglect the auxiliary variables $z_{ij}$ for the bus complex voltage updates.
@@
Our experiments reveal discrepancies between the PyQUBO and DADK implementations, particularly in the number of auxiliary variables generated when encoding problems of identical size. In general, DADK tends to produce more compact formulations with fewer auxiliary variables. We verify that the order reduction preserves the optimality of solutions and maintains consistency in variable assignments corresponding to the minimum-energy configurations.
5 Adiabatic Quantum Algorithm
@@
Iterative schemes are introduced to ensure that AQPF and AQOPF use the unique computational properties of quantum/digital annealers to perform combinatorial PF and OPF, respectively, as presented in Algorithm˜1. Initially, the active power generation vector $\mathbf{P}^{G}$, the demand vectors $\mathbf{P}^{D}$ and $\mathbf{Q}^{D}$, along with the power system admittance matrix $\mathbf{Y}$, are defined (lines 1-4). Next, the voltage real and imaginary parts, $\mathbf{\mu}$ and $\mathbf{\omega}$, as well as their corresponding adjustments, $\Delta{\mu}$ and $\Delta{\omega}$, are initialized by values provided by the user (refer to lines 5–8). These parameters are then used to compute the starting values of the active and reactive power vectors, $\mathbf{P}$ and $\mathbf{Q}$ (see lines 9–10). Note that $P_{1}$ and $Q_{1}$, associated with the slack bus, are excluded to align with the PF problem.
@@
On line 11, the Hamiltonian $H_{\text{obj}}(\cdot)$ is evaluated at the binary vector $\mathbf{x}=[x_{0,0}^{\mu},\dots,x_{n,1}^{\omega}]\in\{0\}^{4n}$ (line 11) using the initial voltage estimations to propose a candidate solution. In line 12, the residual limit, defined as $\epsilon$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$, is initialized, followed by resetting the iteration counter ‘it’ to zero before entering the iterative loop at line 13. After solving the minimization problem ˜15 (line 15), the binary solution vector $\mathbf{x}=[x_{0,0}^{\mu},\dots,x_{n,1}^{\omega}]\in\{0,1\}^{4n}$ derived from the QUBO formulation is employed to revise $\mu_{i}$ and $\omega_{i}$ as specified in line 16 by ˜12. It is important to note that during the voltage update process, any auxiliary variables introduced through the quadratic reformulation of higher-order terms are intentionally omitted. When the computed $H_{\text{obj}}(\mathbf{x})$ falls below the user-defined threshold $\epsilon$, the resulting complex voltages $\mathbf{\mu}+j\mathbf{\omega}$ are considered as the final solution. If not, the base values, $\mu_{i}^{0}$ and $\omega_{i}^{0}$, are reassigned from the current $\mu_{i}$ and $\omega_{i}$ (line 20), and the optimization routine for ˜15 is repeated until convergence is achieved.
@@
For AQOPF, Algorithm˜1 should be modified to align with the OPF problem principle. First, line 1 is not needed, as the generator set points are determined by solving the OPF problem. Moreover, the active and reactive power vectors, $\mathbf{P}$ and $\mathbf{Q}$, are computed only for load buses in lines 9, 10, 17, and 18. Finally, the problem Hamiltonian in lines 11 and 18 should be replaced with ˜29.
@@
9:Calculate $\mathbf{P}=[P_{2},P_{3},\dots,P_{N}]$ using ˜13
10:Calculate $\mathbf{Q}=[Q_{2},Q_{3},\dots,Q_{N}]$ using ˜14
11:Calculate Hamiltonian $H_{\text{obj}}(\mathbf{x})$ using ˜15
@@
16:  Update $\mathbf{\mu}$ and $\mathbf{\omega}$ using ˜12
17:  Recalculate $\mathbf{P}$ using ˜13
18:  Recalculate $\mathbf{Q}$ using ˜14
19:  Recalculate $H_{\text{obj}}(\mathbf{x})$ using ˜15
@@
5.1 Delta Update
The iterative method may exhibit slow convergence when $\Delta\mu_{i}$ and $\Delta\omega_{i}$ are set too low, or if $\mu_{i}^{0}$ and $\omega_{i}^{0}$ are significantly different from the values that fulfill ˜15 and 29. A possible remedy is to employ an adaptive strategy for $\Delta\mu_{i}$ and $\Delta\omega_{i}$, dynamically adjusting their values at each iteration and for each bus $i$ to enhance convergence. In this perspective, the values for $\Delta\mu_{i}$ and $\Delta\omega_{i}$ are a function of the iteration counter ‘it’, as given in:
@@ 5.1 Delta Update
where $\underline{\Delta\mu_{i}}=5\times 10^{-4}$ and $\overline{\Delta\mu_{i}}=4\times 10^{-2}$ are lower and upper bounds of $\Delta\mu_{i}$, respectively; Similarly, $\underline{\Delta\omega_{i}}=1\times 10^{-4}$ and $\overline{\Delta\omega_{i}}=2\times 10^{-2}$ bound $\Delta\omega_{i}$; $e$ is the exponential function; r=0.05r=-0.05 is the rate of decay. In each iteration, $\Delta\mathbf{\mu}$ and $\Delta\mathbf{\omega}$ are updated adaptively and individually; we save two sets of binary variables $\mathbf{x}^{(\text{it}-1)}$ and $\mathbf{x}^{(\text{it}-2)}\in\{0,1\}^{4n}$, which represent the bitstrings for the previous and second previous iterations, respectively. We then use these bitstrings to update $\Delta\mu_{i}$ and $\Delta\omega_{i}$ for each bus $i$. For example, $x_{i,0}^{\mu(\text{it}-2)}=1\land x_{i,1}^{\mu(\text{it}-2)}=0$, $x_{i,0}^{\mu(\text{it}-1)}=1\,\land\,x_{i,1}^{\mu(\text{it}-1)}=0$, and $x_{i,0}^{\mu}=0\land x_{i,1}^{\mu}=1$ indicates that at iteration $\text{it}-2$, the value of $\mu_{i}$ increased, followed by an increase at iteration $\text{it}-1$ and then a decrease at the current iteration ‘it’. This pattern suggests that to prevent $\mu_{i}$ from fluctuations and promote convergence, we should reduce $\Delta\mu_{i}$. Details are given in Algorithm˜2. We employ a similar approach to update $\Delta\mathbf{\omega}$.
@@
6 Results
Four annealers are employed: QA, HA, DAv3, and QIIO. The number of readout samples per execution is fixed across all solvers, ranging from 2,000 to 100,000, and determined by the test case size. A 10-second time limit is assigned to HA. Minor embedding is employed for QA, with the chain strength selected to minimize chain breaks while preserving the relative importance of the original QUBO coefficients. The embedding target graph exhibits sparsity, mirroring the inherently low connectivity structure typical of power systems. For DAv3 and QIIO, the runtime configuration includes a 10-second time limit, 64-bit scaling precision, and a total optimization timeout of 3,600 seconds. Table˜2 summarizes the capability of the solvers to converge to a solution within the residual tolerance of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$. For QA and HA, convergence indicates that an embedding has been found and that the solution is within the accepted residual tolerance. For DAv3 and QIIO, convergence implies that the solution is obtained within the same tolerance.
According to Table˜2, QIIO is capable of handling all test case sizes, including the 1354-bus test case. DAv3, on the other hand, is limited to test cases up to the 89-bus test case due to the constraints on the number of available binary variables. QA and HA are even more restricted and apply only to small test cases. The difference between DAv3 and QIIO lies in their hardware configurations: DAv3 provides a fully connected block of approximately 8,192 binary variables, while QIIO offers multiple fully connected blocks capable of managing up to 100,000 binary variables. This expanded capacity makes QIIO suitable for large-scale problems. Therefore, all further experiments in this paper are conducted using QIIO. Note that the experiments include test cases ranging from 4- to 1354-bus test cases from the Power System Test Cases library [^4] https://pandapower.readthedocs.io/en/v2.2.2. The results for the 4-, 9-, and 14-bus test cases are presented in kaseb2024power , and the results for 5-, 6-, and 15-bus test cases are presented in kaseb2025combinatorial .
Table 2: Performance comparison of different solvers under a user-defined residual of $1\times 10^{-2}$ ${\text{MW}^{2}+\text{MVAR}^{2}}\over 2$.